given f(x)=x-6/5 and g(x)=5x+6, find (g*f)(x) show work
Answer by amarjeeth123(569)

f(x)=x-6/5
g(x)=5x+6
gof(x)=g(f(x))=g(x-6/5)=5(x-6/5)+6=x-6+6=x
Answer=x